# TropNNC: Structured Neural Network Compression Using Tropical Geometry

### v2

@@ Abstract
We present TropNNC, a framework for compressing neural networks with linear and convolutional layers and ReLU activations. TropNNC is a structured compression framework based on a geometrical approach to machine/deep learning, using tropical geometry and extending the work of Misiakos et al. (2022). We use the Hausdorff distance of zonotopes in its standard continuous form to achieve a tighter approximation bound for tropical polynomials compared to previous work. This enhancement leads to the development of an effective compression algorithm that achieves superior functional approximations of neural networks. Our method is significantly easier to implement compared to other frameworks, and does not depend on the availability of training data samples. We validate our framework through extensive empirical evaluations on the MNIST, CIFAR, and ImageNet datasets. Our results demonstrate that TropNNC achieves performance on par with state-of-the-art methods like ThiNet (even surpassing it in compressing linear layers) and CUP. To the best of our knowledge, it is the first method that achieves this using tropical geometry.
Keywords: tropical geometry, zonotopes, Hausdorff distance, neural network compression
@@ 1 Introduction
In recent years, deep neural networks (DNNs) have significantly advanced the field of computer vision, excelling in tasks such as image classification (Krizhevsky et al., 2012; Simonyan and Zisserman, 2015; He et al., 2016; Rawat and Wang, 2017), object detection (Zou et al., 2023), semantic segmentation (Noh et al., 2015; Guo et al., 2018), and image captioning (Jia et al., 2015; Hossain et al., 2019). These networks have outperformed traditional methods reliant on manually crafted visual features. However, despite their success, deploying DNNs on resource-constrained devices, such as mobile phones or embedded systems, remains challenging due to their substantial computational and storage demands. For example, the VGG-16 model (Simonyan and Zisserman, 2015), with its approximately $138$ million parameters, requires over $500$MB of storage and performs about $31$ billion FLOPs to classify a single image. Such resource-intensive models exceed the capabilities of small devices and pose significant challenges, including energy limitations recently highlighted by the deployment of large language models (Samsi et al., 2023). This has led to a growing focus on network compression techniques (Blalock et al., 2020) as a critical area of research.
Initial attempts to reduce network complexity date back to the 1990s, when LeCun et al. (1989) demonstrated that removing negligible weights from a network had minimal impact on performance. This approach, known as neural network pruning, has since evolved. Han et al. (2015) proposed a simple yet effective pruning strategy: iteratively removing connections with weights below a certain threshold and fine-tuning the network to recover accuracy, resulting in highly sparse models.
These methods belong to the category of unstructured pruning, which alters the network’s structure by eliminating individual weights. While effective in reducing network size, unstructured pruning presents challenges in practical applications. Such pruned networks often require specialized hardware and software for efficient inference due to issues like poor cache locality and jumping memory accesses, which can negate the benefits of sparsity and sometimes even degrade performance (Wen et al., 2016).
@@
To overcome the limitations of unstructured pruning, structured pruning methods, such as channel-level pruning, have been proposed (He and Xiao, 2024). Notably, the ThiNet framework by Luo et al. (2017) prunes entire filters or channels, maintaining the network’s original structure. In another work, Wen et al. (2016) proposed the structured sparsity learning approach, which also includes filter- or channel-wise pruning. This structured approach ensures compatibility with existing deep learning libraries and offers several advantages: it significantly reduces memory footprint, facilitates further compression and acceleration through methods like parameter quantization (Gong et al., 2014; Chen et al., 2015; Wu et al., 2016), and accelerates various vision tasks.
Parallel to these advancements, tropical geometry (Maclagan and Sturmfels, 2021) has emerged as a promising mathematical framework with applications in neural networks and machine learning (Maragos et al., 2021). Tropical geometry, rooted in algebraic geometry, operates over the tropical semiring (Cuninghame-Green, 1994), where conventional addition and multiplication are replaced by min and plus operations, respectively. This transformation turns polynomials into piecewise linear functions, making tropical geometry highly relevant to the study of neural networks with piecewise linear activations like ReLU.
@@
Tropical mathematics encompasses a wide range of applications, which include operations research and scheduling (Cuninghame-Green, 1979), control and optimization (Baccelli et al., 2001; Cohen et al., 2004; Butkovič, 2010; Akian et al., 2012), mathematical physics (Litvinov et al., 2001), speech recognition (Mohri et al., 2002), polyhedral geometry (Gaubert and Katz, 2011), tropical geometry (Maclagan and Sturmfels, 2021), dynamical systems on weighted lattices (Maragos, 2017), sparsity theory on weighted lattices (Tsilivis et al., 2022), matrix factorization (Kordonis et al., 2024), finite state transducers (Theodosis and Maragos, 2018), convex regression (Maragos and Theodosis, 2020), and machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008). Recently, tropical geometry has been applied to the theoretical study of neural networks. For example, Zhang et al. (2018) demonstrated the equivalence of ReLU-activated neural networks with tropical rational mappings, and used zonotopes to compute a bound on the number of linear regions of networks, equal to the one in (Montufar et al., 2014). Other works, like those of Charisopoulos and Maragos (2018); Alfarra et al. (2023); Smyrnis and Maragos (2020), have used tropical geometry to compute bounds on the number of linear regions in neural networks and to develop methods for pruning and compressing these networks.
@@
In this paper, we explore the application of tropical geometry to neural network compression. We build on the work of Misiakos et al. (2022), who utilized tropical geometry for structured neural network compression by geometrically approximating the zonotopes corresponding to the network layers. Our contributions include:
  • Obtaining a tighter bound for the functional approximation of tropical polynomials than that presented in (Misiakos et al., 2022) using the standard, continuous Hausdorff distance.
  • Designing an algorithm that leverages tropical geometry and the Hausdorff distance for the structured compression of neural networks. Our algorithm refines the choice of weights in the compressed network from the clustering done in (Misiakos et al., 2022) and we also provide an iterative technique to further optimize them. Our method is simple to implement and does not depend on the availability of training data samples like other frameworks such as ThiNet (Luo et al., 2017). It is also the first method based on the theory of tropical geometry that compresses convolutional layers.
@@
  • Evaluating our algorithm through experiments on MNIST, Fashion-MNIST, CIFAR-10, CIFAR-100, and ImageNet datasets. Our method outperforms that of Misiakos et al. (2022) and achieves competitive or superior performance compared to ThiNet, especially in compressing linear layers, and superior performance compared to CUP (Duggal et al., 2021), particularly in the VGG architecture.
This work demonstrates the potential of tropical geometry in enhancing neural network compression techniques, providing a theoretical foundation and a practical algorithm that improves efficiency and maintains accuracy.
2 Related Work
Seminal works have showcased that neural networks tend to be hyperparameterized. Denil et al. (2013) demonstrated that neural networks can be reconstructed using a small subset of parameters, which sparked significant interest in network compression techniques. Blalock et al. (2020) reviewed various methods for reducing model size without compromising performance, highlighting the ongoing advancements in this area.
Early methods for network compression focused on unstructured pruning. LeCun et al. (1989) showed that removing unimportant weights leads to minimal performance loss. This approach was further refined by Han et al. (2015), who proposed pruning weights below a certain threshold followed by fine-tuning the network. Despite their effectiveness, these methods require specialized hardware for efficient inference due to their unstructured nature.
Structured pruning techniques maintain the network’s architecture, facilitating better compatibility and efficiency. Wen et al. (2016) introduced structured sparsity learning to compress networks by finding compact representations of filters and channels, preserving the architecture. Luo et al. (2017) proposed ThiNet, a method that greedily prunes filters based on their impact on the next layer, thereby maintaining network structure and efficiency. Yu et al. (2018) presented the Neuron Importance Score Propagation (NISP) algorithm, which extended the ideas of ThiNet by pruning neurons and channels based on their impact on the final layer output. Lin et al. (2020) introduced HRank, which uses feature map information to rank and prune filters. Duggal et al. (2021) proposed Cluster Pruning (CUP), which prunes similar filters by clustering them based on features derived from both incoming and outgoing weight connections. CUP addresses the limitations of prior work by efficiently determining the ideal number of filters to prune in each layer and integrating pruning into the initial training phase, leading to significant savings in training time and maintaining performance. Wang et al. (2021) introduced a novel approach to pruning convolutional neural networks that focuses on reducing structural redundancy rather than merely pruning individual neurons or channels. Their method identifies and eliminates redundant structures within the network architecture, thereby achieving more efficient pruning without compromising the network’s performance. Smyrnis and Maragos (2020) proposed a tropical division algorithm for structured network compression, further enhancing the capabilities of structured pruning techniques. Misiakos et al. (2022) developed Neural Path K-means, which uses clustering techniques based on tropical geometry and functional approximation for structured compression.
Beyond pruning, other state-of-the-art strategies for network compression include parameter quantization and low-rank approximation. Parameter quantization, as explored by Zhou et al. (2017), Zhou et al. (2018) and Han et al. (2015), reduces the number of bits per parameter, significantly lowering memory and computational requirements. Behdin et al. (2023) presented an optimization-based framework for efficient post-training quantization of large language models. Polino et al. (2018) presented quantized distillation, a method that combines quantization with teacher-student distillation. Low-rank approximation (LRA) techniques decompose weight matrices into smaller matrices, reducing parameters and computations. Studies by Denton et al. (2014), Jaderberg et al. (2014), Sindhwani et al. (2015), Ioannou et al. (2016), Tai et al. (2016), and Chen et al. (2018) have shown that LRA can achieve significant speedups and compression with minimal accuracy loss, although it typically requires iterative optimization. These methods can be combined with structured pruning approaches like ThiNet to achieve further model compression and efficiency.
Our framework is a structured pruning method that extends the approach of Misiakos et al. (2022). It prunes networks layer-by-layer, aiming to minimize the error introduced to the input of the next layer. In that sense, it is similar to the ThiNet framework. However, our algorithm is simpler to implement and does not require a training dataset. Our framework leverages clustering and can be naturally extended to non-uniform pruning, drawing parallels to the CUP framework. In fact, it improves upon steps 1, 2, and 3 of CUP. For step 1, we use similar clustering vectors (referred to as filter features by CUP authors Duggal et al. (2021)) for linear layers, and introduce a new method for constructing clustering vectors in convolutional layers. For step 2, we introduce two approaches for choosing the distance threshold of the hierarchical clustering. Our main contribution lies in step 3, where we employ tropical geometry to select a better representative for each cluster, rather than simply choosing the neuron or channel with the most energy. While it is tempting to compare our framework with those of NISP (Yu et al., 2018) and HRank (Lin et al., 2020) given that they are both structured pruning methods, it is important to recognize the distinctions between them and ours. NISP and HRank are advanced frameworks specifically designed to address the limitations of layer-by-layer pruning approaches, such as those employed by ThiNet and our own method. In contrast, our framework leverages a rich theoretical foundation and does not rely on training data samples, offering a unique advantage in certain contexts. We believe our results are promising and present significant advantages. Our work highlights the potential of tropical geometrical techniques and paves the way for further research that could enhance their competitiveness with current state-of-the-art methods.
@@ 3 Tropical Geometry of Neural Networks
In this section, we introduce the fundamental background of tropical geometry and its applications to neural network compression.
3.1 Background on Tropical Geometry
@@
Tropical algebra is a field of mathematics that studies the tropical semiring. Tropical geometry is the counterpart of algebraic geometry in the tropical setting. The tropical semiring can refer to either the min-plus semiring or the max-plus semiring (Butkovič, 2010). In this work, we adhere to the convention of using the max-plus semiring $(\mathbb{R}_{\mathrm{max}},\vee,+)$, defined as the set $\mathbb{R}_{\mathrm{max}}=\mathbb{R}\cup\{-\infty\}$ equipped with two binary operations: $\vee$ (ordinary max) and $+$ (ordinary sum).
3.1.1 Tropical Polynomials and Rational Functions
where $[n]:=\{1,...,n\}$. Here, $n$ represents the rank of the tropical polynomial. Each monomial term $\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\}$ of the polynomial has an exponent or slope $\mathbf{a}_{i}\in\mathbb{R}^{d}$ and a coefficient or bias $b_{i}\in\mathbb{R}$. Each monomial term corresponds to a plane in $\mathbb{R}^{d+1}$. Consequently, tropical polynomials are piecewise linear convex functions. Specifically, every tropical polynomial is a continuous piecewise linear convex function, and every continuous piecewise linear convex function can be expressed (though not uniquely) as a tropical polynomial (Maclagan and Sturmfels, 2021). The set of tropical polynomials in $\mathbf{x}$ defines the semiring $\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$. Figures 1(a), 1(b) illustrate examples of tropical polynomials in one and multiple variables, respectively.
Tropical rational functions extend the concept of rational polynomial functions to tropical algebra. A tropical rational function is defined as the tropical multiplication of a tropical polynomial $p$ by the tropical multiplicative inverse $q^{-1}$ of another tropical polynomial $q$. In conventional arithmetic, this operation corresponds to the difference of two tropical polynomials $p$ and $q$:
$r(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x})$
Tropical rational functions correspond to general piecewise linear functions. Specifically, every tropical rational function is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational function. Figures 1(c), 1(d) provide examples of tropical rational functions.
(a) $f(x)=\max\{x,1\}$
(b) $f(x,y)=\max\{x,y,1\}$
(c) $f(x)=\max\{2x-2,-x\}-\max\{-2x-2,x\}$
(d) $f(x,y)=\max\{x,-x\}-\max\{y,-y\}$
3.1.2 Newton Polytopes
As with algebraic geometry (Hartshorne, 2013), in tropical geometry we can define the Newton polytope (Monical et al., 2019) of a tropical polynomial. Newton polytopes can be used to analyse the behavior of a polynomial. They connect tropical geometry with polytope theory, an extensively studied field of mathematics (Ziegler, 2012). For a tropical polynomial $f$ as defined in (1), we define its Newton polytope as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$
$\mathrm{Newt}(f):=\mathrm{conv}\left\{\mathbf{a}_{i},i\in[n]\right\}.$
The following proposition allows us to calculate the (extended) Newton polytope of expressions of tropical polynomials.
Proposition 1 (Zhang et al., 2018)
$\mathrm{ENewt}(f+g)$ $=\mathrm{ENewt}(f)\oplus\mathrm{ENewt}(g).$
Here, $\vee$ denotes the tropical addition (taking the maximum), and $+$ denotes the tropical multiplication (ordinary addition). The Minkowski sum $\oplus$ of two polytopes (or more generally subsets of $\mathbb{R}^{d}$) $P$ and $Q$ is defined as:
$P\oplus Q:=\{p+q\mid p\in P,q\in Q\}.$
The extended Newton polytope provides a geometrical interpretation for studying tropical polynomials. The following propositions establish the relationship between tropical polynomials and their extended Newton polytopes. The upper envelope or upper hull $UF(P)$ of an extended Newton polytope is defined as the set of all points $(\mathbf{a}^{T},b)$ of the polytope $P$ that are not ”shadowed” by any other part of the polytope when viewed from above (last dimension). This means that there is no $b^{\prime}>b$ such that $(\mathbf{a}^{T},b^{\prime})$ belongs to $P$. We have the following useful lemma (proof in Appendix A.1).
Lemma 2
An immediate consequence of the above lemma is the following theorem:
Theorem 3
Let $f\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be a tropical polynomial. The values of $f$ are fully determined by the upper envelope $UF(\mathrm{ENewt}(f))$ of its extended Newton polytope.
Theorem 4 (Charisopoulos and Maragos, 2018)
The linear regions of a tropical polynomial $f\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ are in one-to-one correspondence with the vertices of $UF(\mathrm{ENewt}(f))$. In fact, each vertex $(\mathbf{a}_{i}^{T},b_{i})\in UF(\mathrm{ENewt}(f))$ gives exactly one linear region $D$ where $f(\mathbf{x})=\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}$.
The above theorem highlights the direct relationship between the geometric structure of the extended Newton polytope and the piecewise linear nature of the tropical polynomial. An immideate consequence of Theorems 3 and 4 is the following theorem:
Theorem 5
$f=g\Leftrightarrow UF(\mathrm{ENewt}(f))=UF(\mathrm{ENewt}(g))$
Example 1
Consider the polynomials corresponding to Figure 2.
$f(x,y)=\max\{0,-y+1,y+1\}$
$g(x,y)=\max\{x+1,-x+1\}.$
We have that
$(f\vee g)(x,y)=\max\{0,x+1,y+1,-y+1,-x+1\}$
$(f+g)(x,y)=\max\{x+1,-x+1,x-y+2,-x-y+2,x+y+2,-x+y+2\}.$
$(f\vee g)(x,y)=\max\{x+1,y+1,-y+1,-x+1\},$
(a) $\mathrm{ENewt}(f)$
(b) $\mathrm{ENewt}(g)$
(c) $\mathrm{ENewt}(f\vee g)$
(d) $\mathrm{ENewt}(f+g)$
3.1.3 Zonotopes
$Z:=\bigoplus_{i\in[n]}g_{i}$
The line segments $g_{i}$ are referred to as the zonotope’s generators. Each generator contributes to the shape and size of the zonotope, which is formed by the space swept out by these line segments.
Alternatively, a zonotope can be expressed equivalently by a set of vectors $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$ and a starting point $\mathbf{s}\in\mathbb{R}^{d}$. By taking the generators to be the segments $[\mathbf{0},\mathbf{v}_{1}],\ldots,[\mathbf{0},\mathbf{v}_{n}]$ and translating the first segment by $\mathbf{s}$, we obtain the equivalent form:
In this context, the vectors $\mathbf{v}_{i}$ are sometimes referred to as the zonotope’s generators, meaning the segments $[\mathbf{0},\mathbf{v}_{i}]$. When the starting point $\mathbf{s}$ is not mentioned, it is assumed to be the origin $\mathbf{0}$.
$\mathbf{u}=\mathbf{s}+\sum_{i\in I}\mathbf{v}_{i},$
where $I\subseteq[n]$. It can be shown that zonotopes have up to exponential many verticesIn fact, for generators in general positions they have exactly $2\sum_{j=0}^{d-1}nCr(n-1,j)$ vertices (Gritzmann and Sturmfels, 1993).
Zonotopes exhibit several interesting properties. Notably, they are centrally symmetric. This symmetry, along with their convex structure, make zonotopes particularly useful in various fields, including optimization (Bern and Eppstein, 2001) and computational geometry (Ziegler, 2012).
@@ 3.2 Neural Networks with Piecewise Linear Activations
Tropical geometry provides a mathematical framework for analyzing neural networks with piecewise linear activation functions. Leveraging the properties of tropical polynomials and Newton polytopes, it is used to model the behavior of various activation functions commonly used in neural networks. In this work, we focus on ReLU-activated networks. In this section, we outline the basic tropical geometrical properties of ReLU-activated networks and establish our notations.
ReLU Activations. Consider a network which consists of an input layer $\mathbf{x}=(x_{1},\ldots,x_{d})$, a hidden layer $\mathbf{f}=(f_{1},\ldots,f_{n})$ of ReLU units, and an output layer $\mathbf{v}=(v_{1},\ldots,v_{m})$. The input, hidden, and output layers are connected through linear transformations represented by matrices $\mathbf{A}$ and $\mathbf{C}$. Each neuron $i$ has input weights and bias given by $\mathbf{A}_{i,:}=(\mathbf{a}_{i}^{T},b_{i})$ and output weights $\mathbf{C}_{:,i}^{T}=(c_{1i},\ldots,c_{mi})$. We assume the output layer has no bias. Such a network is depicted in Figure 3.
Figure 3: Neural network with one hidden ReLU layer. The first linear layer has weights $\{\mathbf{a}_{i}^{T}\}$ with bias $\{b_{i}\}$ corresponding to $i$-th node $\forall i\in[n]$ and the second has weights $\{c_{ji}\}$, $\forall j\in[m],i\in[n]$
@@
This expression represents a tropical polynomial of rank 2, with one term being the constant 00. The extended Newton polytope $\mathrm{ENewt}(f_{i})$ of $f_{i}$ is an edge with one endpoint at the origin $\mathbf{0}$ and the other endpoint at $(\mathbf{a}_{i}^{T},b_{i})$. The $j$-th component of the output layer $v_{j}$ can be computed as follows:
@@
In the above expression, $|c_{ji}|f_{i}$ are tropical polynomials. Thus, $p_{j}$ and $q_{j}$ are tropical polynomials formed by the addition of tropical polynomials. Consequently, $v_{j}$ is a tropical rational function. We call $p_{j}$ the positive polynomial and $q_{j}$ the negative polynomial of $v_{j}$. This result can be extended to deeper networks, as suggested by the following proposition:
Theorem 6 (Zhang et al., 2018)
A ReLU-activated deep neural network $\mathbf{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is a tropical rational mapping (vector whose elements are tropical rational functions).

### v3

@@ Abstract
We present TropNNC, a framework for compressing neural networks with linear and convolutional layers and ReLU activations using tropical geometry. By representing a network’s output as a tropical rational function, TropNNC enables structured compression via reduction of the corresponding tropical polynomials. Our method refines the geometric approximation of previous work by adaptively selecting the weights of retained neurons. Key contributions include the first application of tropical geometry to convolutional layers and the tightest known theoretical compression bound. TropNNC requires only access to network weights – no training data – and achieves competitive performance on MNIST, CIFAR, and ImageNet, matching strong baselines such as ThiNet and CUP.
11footnotetext: Corresponding author.
@@ 1 Introduction
Deploying deep neural networks on resource-constrained devices remains challenging due to their substantial computational and storage demands. Initial attempts to reduce network complexity include unstructured pruning. which alters the network’s structure by eliminating individual weights. While effective in reducing network size, unstructured pruning presents challenges in practical applications (Wen et al., 2016). To overcome its limitations, structured pruning methods, such as channel-level pruning, have been proposed (He and Xiao, 2024). Notably, the ThiNet framework by Luo et al. (2017) prunes entire filters or channels, maintaining the network’s original structure. This structured approach ensures compatibility with existing deep learning libraries and offers several advantages: it significantly reduces memory footprint, and facilitates further compression and acceleration through methods like parameter quantization.
@@
Parallel to these advancements, tropical geometry (Maclagan and Sturmfels, 2021) has emerged as a promising mathematical framework with applications in machine learning (Maragos et al., 2021; Gärtner and Jaggi, 2008) and beyond. Recently, it has been applied to the theoretical study of neural networks. For example, Zhang et al. (2018) demonstrated the equivalence of ReLU-activated neural networks with tropical rational mappings. Together with other works, like those of Charisopoulos and Maragos (2018); Alfarra et al. (2023), they used tropical geometry to compute bounds on the number of linear regions of neural networks equal to the one in (Montufar et al., 2014). Smyrnis and Maragos (2020) use tropical geometry to develop pruning methods.
Contributions.
@@
In this paper, we explore the application of tropical geometry in the compression of ReLU neural networks. Our contributions include:
  • Proposing TropNNC, an algorithm that leverages tropical geometry and the Hausdorff distance for the structured compression of neural networks. TropNNC compresses convolutional networks layer-wise by approximating the zonotopes corresponding to each layer using an iterative process. Our algorithm is data agnostic: it does not require a training dataset.
  • Refining the bound for the functional approximation of tropical polynomials presented in (Misiakos et al., 2022), and using it to provide stronger theoretical compression guarantees.
@@
  • Evaluating our algorithm empirically on MNIST, CIFAR and ImageNet datasets. Our method outperforms prior tropical geometrical pruning methods, and achieves competitive performance compared to the data-driven ThiNet, and superior performance compared to CUP (Duggal et al., 2021), particularly in the VGG architecture.
This work demonstrates the potential of tropical geometry in enhancing neural network compression techniques. To the best of our knowledge, it is the first tropical geometric pruning method that compresses convolutional layers and is competitive with strong baselines. Even if currently not state-of-the-art, we believe that our contribution is a necessary step towards this direction. Related work is provided after the definition of our algorithm to enable detailed comparisons. Proofs of theoretical results are provided in the appendix.
@@ 2 Tropical Geometry of Neural Networks
Tropical algebra studies matrix-vector operations based on the arithmetic of the tropical semiring Cuninghame-Green (1979); Butkovič (2010). Tropical geometry is the counterpart of algebraic geometry in the tropical setting. The tropical semiring can refer to either the min-plus semiring or the max-plus semiring. In this work, we adhere to the convention of using the max-plus semiring $(\mathbb{R}_{\mathrm{max}},\vee,+)$, defined as the set $\mathbb{R}_{\mathrm{max}}=\mathbb{R}\cup\{-\infty\}$ equipped with two binary operations: $\vee$ (ordinary max) and ++ (ordinary sum).
Within the max-plus semiring, we can define tropical polynomials, which correspond to convex piecewise linear functions, and tropical rational functions, which correspond to general continuous piecewise linear functions. We can also define Newton polytopes of tropical polynomials, which connect tropical algebra with polytope theory. For a detailed introduction to these concepts refer to the appendix.
@@ 2.1 Neural Networks with Piecewise Linear Activations
Tropical geometry provides a mathematical framework for analyzing neural networks with piecewise linear activation functions. In this work, we focus on ReLU-activated networks.
ReLU Activations.
Consider a network which consists of an input layer $\mathbf{x}=(x_{1},\ldots,x_{d})$, a hidden layer $\mathbf{f}=(f_{1},\ldots,f_{n})$ of ReLU units, and an output layer $\mathbf{v}=(v_{1},\ldots,v_{m})$. The input, hidden, and output layers are connected through linear transformations represented by matrices $\mathbf{A}$ and $\mathbf{C}$. Each neuron $i$ has input weights and bias given by $\mathbf{A}_{i,:}=(\mathbf{a}_{i}^{T},b_{i})$ and output weights $\mathbf{C}_{:,i}^{T}=(c_{1i},\ldots,c_{mi})$. We assume the output layer has no bias. Such a network is depicted in Figure 1.
Figure 1: Neural network with one hidden ReLU layer. The first linear layer has weights $\{\mathbf{a}_{i}^{T}\}$ with bias $\{b_{i}\}$ corresponding to node $i\in[n]$ and the second has weights $\{c_{ji}\}$ between nodes $j\in[m],i\in[n]$
@@
This expression represents a tropical polynomial of rank 2, with one term being the constant 0. The extended Newton polytope $\mathrm{ENewt}(f_{i})$ of $f_{i}$ is an edge with one endpoint at the origin $\mathbf{0}$ and the other endpoint at $(\mathbf{a}_{i}^{T},b_{i})$. The $j$-th component of the output layer $v_{j}$ can be computed as follows:
@@
In the above expression, $|c_{ji}|f_{i}$ are tropical polynomials. Thus, $p_{j}$ and $q_{j}$ are tropical polynomials formed by the addition of tropical polynomials. Consequently, $v_{j}$ is a tropical rational function. We call $p_{j}$ the positive polynomial and $q_{j}$ the negative polynomial of $v_{j}$.
Zonotopes.